Ryan has 13 coins worth $1.09. If the coins are all quarters and pennies, how many quarters and pennies does he have?

Answer by stanbon(75887)

Ryan has 13 coins worth $1.09. If the coins are all quarters and pennies, how many quarters and pennies does he have?
-----------------------
q + p = 13 coins
25q + p = 109 cents
---------------------------------
Subtract and solve for "q":
24q = 96
----
q = 4 (# of quarters)
----
p = 13-q = 9 (# of pennies)
------------
Cheers,
Stan H.
-----------------